# MnLargeSymbols’164 MnLargeSymbols’171

### v1

@@ 16.4 Duality in syntomic cohomology
Assume that $\mathcal{X}$ is proper of relative dimension $d$. As in §16.4, we will write $\mathbf{Syn}^{c}$ rather than $\mathbf{Syn}^{c\text{-}\infty}$ when considering the analytic syntomic cohomology. Take $\mathcal{X}_{3}\rightarrow\mathcal{X}_{1}\times\mathcal{X}_{2}$ to be the diagonal of $\mathcal{X}=\mathcal{X}_{i}$ and $\mathbb{E}_{3}=\left(\mathcal{O},\mathrm{Fil}^{\bullet}\mathcal{O},d,\mathcal{O}_{\mathfrak{X}},\sigma\right)=:\mathbb{I}$ the unit object. Then we get a pairing (cfr. Remark 15.5) :
@@
$\tau_{\mathrm{rig},U,\ell,c}^{i}\colon\mathrm{H}^{i}\left(\mathbf{Syn}_{\mathrm{rig},U^{o}\subset W_{\ell},\alpha_{k}}^{c}\left(\mathbb{E}^{\vee},P_{2},r_{2}\right)\right)\longrightarrow\mathrm{H}^{i}\left(\mathbf{Syn}^{c\text{-}\infty}\left(\mathbb{E}^{\vee},P_{2},r_{2}\right)\right).$
@@
Assume $\left(P_{1}\ast P_{2},r_{1}+r_{2}\right)$ is admissible. Let $\in\left\{\emptyset,c\text{-}\infty\right\}$. Then, for $f\in\mathrm{H}^{i}\left(\mathbf{Syn}^{?}\left(\mathbb{E},P_{1},r_{1}\right)\right)$ and $g\in\mathrm{H}^{2d+1-i}\left(\mathbf{Syn}_{\mathrm{rig},U\subset W_{\ell},\alpha_{k}}^{c}\left(\mathbb{E}^{\vee},P_{2},r_{2}\right)\right)$, we have
@@ 16.5 Dualities for coefficients with good reduction
Assume that $W_{\ell}$ is smooth and let $U\subset W_{\ell}$ be an open dense subscheme. Assume that $\mathbb{E}=(\mathcal{E},\mathfrak{E})$ satisfies the Assumption of §6.4. It follows from [EY, Prop. 3.15] that $\mathrm{H}^{i}_{\rm rig,c}\bigl((U^{o}\subset W_{\ell},\alpha_{k}^{\rm hor})/K_{0},\mathfrak{E}^{\rm conv}\bigr)$ and $\mathrm{H}^{i}_{\rm rig,c}\bigl((U^{o}\subset W_{\ell},\alpha_{k}^{\rm hor})/K,\mathfrak{E}^{\rm conv}\bigr)$ coincides with Berthelot’s rigid cohomology $\mathrm{H}^{i}_{{\rm rig},c}\bigl(U^{o}/K_{0}\bigr)$ and $\mathrm{H}^{i}_{{\rm rig},c}\bigl(U^{o}/K\bigr)$ respectively. In particular, these groups vanish for $i\geq 2d+1$ and we have trace maps
@@
For $\left(P,r\right)$ admissible, $P(\Phi_{L})$ is an isomorphism on $\mathrm{H}^{2d}_{{\rm rig},c}\bigl(U^{o}/K_{0})$ and we get an analogue of Lemma 15.2 for $\mathbb{E}=\mathbb{I}=(\mathcal{O},\mathfrak{O})$ the unit object, namely
@@
For every element $f\in\mathrm{H}^{i}\left(\mathbf{Syn}_{\mathrm{rig},U\subset W_{l},\alpha_{k}^{hor}}^{?}\left(\mathbb{E},P_{1},r_{1}\right)\right)$ with $\in\left\{\emptyset,c\text{-}\infty\right\}$ and for every element $g\in\mathrm{H}^{2d+1-i}\left(\mathbf{Syn}_{\mathrm{rig},U^{o}\subset W_{\ell},\alpha_{k}^{\rm hor}}^{c}\left(\mathbb{E}^{\vee},P_{2},r_{2}\right)\right)$, we have
@@
For $f\in\mathrm{H}^{i}\left(\mathbf{Syn}^{?}\left(\mathbb{E},P_{1},r_{1}\right)\right)$ with $\in\left\{\emptyset,c\text{-}\infty\right\}$ and $g$ as in Lemma 16.6, we have
@@ 16.6 Adjointness of Gysin morphisms and pull-back morphisms
We suppose to be in the setting of §13, except we replace the complexes $\mathbf{Syn}\left(\mathcal{E},\mathfrak{E},r\right)$ with the syntomic polynomial complexes $\mathbf{Syn}\left(\mathbb{E},P,r\right)$ of §14, with and without support conditions. We use the abbreviation $\mathbf{Syn}^{c}$ for $\mathbf{Syn}^{c\text{-}\infty}$. In particular, there is a closed immersion $\iota$ of $\mathcal{F}$ in $\mathcal{X}$.
@@
$\mathrm{Gys}_{\mathrm{syn},F}^{P,r}\colon\mathrm{H}^{i-2}\left(\mathbf{Syn}\left(\mathbb{E}_{\mid F}\left(-1\right),P,r\right)\right)\longrightarrow\mathrm{H}^{i}\left(\mathbf{Syn}\left(\mathbb{E},P,r\right)\right)$
@@
The map $\mathrm{Gys}_{\mathrm{syn},F}^{\mathrm{top},P,r}$ is compatible with the trace maps on $X$ and $F$ respectively. Moreover,
@@
for $f\in\mathrm{H}^{i-2}\left(\mathbf{Syn}\left(\mathbb{E}_{\mid F}\left(-1\right),P_{1},r_{1}\right)\right)$ and $g\in\mathrm{H}^{2d+1-i}\left(\mathbf{Syn}^{c}\left(\mathbb{E}^{\vee},P_{2},r_{2}\right)\right)$ and
@@
The compatibility of $\mathrm{Gys}_{\mathrm{syn}^{c},F}^{\mathrm{top},P,r}$ with the trace maps reduces to an analogous statement on de Rham cohomology (in view of Definition 14) and this is proven in [LLZ, Prop. 4.3.7]. We prove the first equality. We have to show the compatibility of the pairings with values in $\mathrm{H}^{2d+1}\left(\mathbf{Syn}^{c}\left(\mathbb{I},P,r\right)\right)$ and $\mathrm{H}^{2d-1}\left(\mathbf{Syn}^{c}\left(\mathbb{I}_{F}\left(-1\right),P,r\right)\right)$ respectively via $\mathrm{Gys}_{\mathrm{syn}^{c},F}^{\mathrm{top},P,r}$, namely that we have

### v2

@@ 16.4 Duality in syntomic cohomology
Assume that $\mathcal{X}$ is proper of relative dimension $d$. As in §16.1, we will write $\mathbf{Syn}^{c}$ rather than $\mathbf{Syn}^{c\text{-}\infty}$ when considering the analytic syntomic cohomology. Take $\mathcal{X}_{3}\rightarrow\mathcal{X}_{1}\times\mathcal{X}_{2}$ to be the diagonal of $\mathcal{X}=\mathcal{X}_{i}$ and $\mathbb{E}_{3}=\left(\mathcal{O},\mathrm{Fil}^{\bullet}\mathcal{O},d,\mathcal{O}_{\mathfrak{X}},\sigma\right)=:\mathbb{I}$ the unit object.
@@
$\tau_{\mathrm{rig},U,\ell,c}^{i}\colon\mathrm{H}^{i}\left(\mathbf{Syn}_{\mathrm{rig},U^{o}\subset W_{\ell},\alpha_{k}}^{c}\left(\mathbb{E}^{\vee},P_{2},r_{2}\right)\right)\longrightarrow\mathrm{H}^{i}\left(\mathbf{Syn}^{c-\infty}\left(\mathbb{E}^{\vee},P_{2},r_{2}\right)\right).$
@@
Assume $\left(P_{1}\ast P_{2},r_{1}+r_{2}\right)$ is admissible. Let $\in\left\{\emptyset,c-\infty\right\}$. Then, for $f\in\mathrm{H}^{i}\left(\mathbf{Syn}^{?}\left(\mathbb{E},P_{1},r_{1}\right)\right)$ and $g\in\mathrm{H}^{2d+1-i}\left(\mathbf{Syn}_{\mathrm{rig},U\subset W_{\ell},\alpha_{k}}^{c}\left(\mathbb{E}^{\vee},P_{2},r_{2}\right)\right)$, we have
@@ 16.5 Dualities for coefficients with good reduction
Assume that $W_{\ell}$ is smooth and let $U\subset W_{\ell}$ be an open dense subscheme. Assume that $\mathbb{E}=(\mathcal{E},\mathfrak{E})$ satisfies the Assumption of §6.4. It follows from [EY20, Prop. 3.15] that $\mathrm{H}^{i}_{\rm rig,c}\bigl((U^{o}\subset W_{\ell},\alpha_{k}^{\rm hor})/K_{0},\mathfrak{E}^{\rm conv}\bigr)$ and $\mathrm{H}^{i}_{\rm rig,c}\bigl((U^{o}\subset W_{\ell},\alpha_{k}^{\rm hor})/K,\mathfrak{E}^{\rm conv}\bigr)$ coincides with Berthelot’s rigid cohomology $\mathrm{H}^{i}_{{\rm rig},c}\bigl(U^{o}/K_{0}\bigr)$ and $\mathrm{H}^{i}_{{\rm rig},c}\bigl(U^{o}/K\bigr)$ respectively. In particular, these groups vanish for $i\geq 2d+1$ and we have trace maps
@@
For $\left(P,r\right)$ admissible, $P(\Phi_{L})$ is an isomorphism on $\mathrm{H}^{2d}_{{\rm rig},c}\bigl(U^{o}/K_{0})$ and we get from Lemma 14.2 for $\mathbb{E}=\mathbb{I}=(\mathcal{O},\mathfrak{O})$ the unit object that
@@
For every element $f\in\mathrm{H}^{i}\left(\mathbf{Syn}_{\mathrm{rig},U\subset W_{l},\alpha_{k}^{hor}}^{?}\left(\mathbb{E},P_{1},r_{1}\right)\right)$ with $\in\left\{\emptyset,c-\infty\right\}$ and for every element $g\in\mathrm{H}^{2d+1-i}\left(\mathbf{Syn}_{\mathrm{rig},U^{o}\subset W_{\ell},\alpha_{k}^{\rm hor}}^{c}\left(\mathbb{E}^{\vee},P_{2},r_{2}\right)\right)$, we have
@@
For $f\in\mathrm{H}^{i}\left(\mathbf{Syn}^{?}\left(\mathbb{E},P_{1},r_{1}\right)\right)$ with $\in\left\{\emptyset,c-\infty\right\}$ and $g$ as in Lemma 16.6, we have
@@ 16.6 Adjointness of Gysin morphisms and pull-back morphisms
We suppose to be in the setting of §13, except we replace the complexes $\mathbf{Syn}\left(\mathcal{E},\mathfrak{E},r\right)$ with the syntomic polynomial complexes $\mathbf{Syn}\left(\mathbb{E},P,r\right)$ of §14, with and without support conditions. We use the abbreviation $\mathbf{Syn}^{c}$ for $\mathbf{Syn}^{c-\infty}$. In particular, there is a closed immersion $\iota$ of $\mathcal{F}$ in $\mathcal{X}$.
@@
$\mathrm{Gys}_{\mathrm{syn},F}^{P,r}\colon\mathrm{H}^{i-2}\left(\mathbf{Syn}\left(\mathbb{E}_{\mid F}\left(-1\right),P,r\right)\right)\longrightarrow\mathrm{H}^{i}\left(\mathbf{Syn}\left(\mathbb{E},P,r\right)\right),$
@@
The map $\mathrm{Gys}_{\mathrm{syn},F}^{\mathrm{top},P,r}$ is compatible with the trace maps on $X$ and $F$ respectively and
@@
for $f\in\mathrm{H}^{i-2}\left(\mathbf{Syn}\left(\mathbb{E}_{\mid F}\left(-1\right),P_{1},r_{1}\right)\right)$ and $g\in\mathrm{H}^{2d+1-i}\left(\mathbf{Syn}^{c}\left(\mathbb{E}^{\vee},P_{2},r_{2}\right)\right)$. Analogously,
@@
The compatibility of $\mathrm{Gys}_{\mathrm{syn}^{c},F}^{\mathrm{top},P,r}$ with the trace maps reduces to an analogous statement on de Rham cohomology (in view of Definition 14) and this is proven in [LLZ23, Prop. 4.3.7]. We prove the first equality. We have to show the compatibility of the pairings with values in $\mathrm{H}^{2d+1}\left(\mathbf{Syn}^{c}\left(\mathbb{I},P,r\right)\right)$ and $\mathrm{H}^{2d-1}\left(\mathbf{Syn}^{c}\left(\mathbb{I}_{F}\left(-1\right),P,r\right)\right)$ respectively via $\mathrm{Gys}_{\mathrm{syn}^{c},F}^{\mathrm{top},P,r}$, namely that we have
@@
Références
  • [AI12] F. Andreatta and A. Iovita (2012) Semistable sheaves and comparison isomorphisms in the semistable case. Rend. Semin. Mat. Univ. Padova 128, pp. 131–285 (2013). External Links: ISSN 0041-8994, Document, Link, MathReview (Francesco Sala) Cited by: §10, §11, §11.2, §11.2, §4.1, §7, §7, §7, Remark 7.2, §7, §8.
  • [BER96] P. Berthelot (1996) Cohomologie rigide et cohomologie rigide à supports propres. Prépubication (), pp. . External Links: ISSN , Document, Link Cited by: §2.
  • [BO78] P. Berthelot and A. Ogus (1978) Notes on crystalline cohomology. Princeton University Press, Princeton, NJ; University of Tokyo Press, Tokyo. External Links: ISBN 0-691-08218-9, MathReview (G. Horrocks) Cited by: §5.
  • [BER86] P. Berthelot (1986) Géométrie rigide et cohomologie des variétés algébriques de caractéristique $p$. pp. 3, 7–32. Note: Introductions aux cohomologies $p$-adiques (Luminy, 1984) External Links: ISSN 0037-9484, MathReview (F. Baldassarri) Cited by: §6.1, §6.3.
  • [BES00] A. Besser (2000) A generalization of Coleman’s $p$-adic integration theory. Invent. Math. 142 (2), pp. 397–434. External Links: ISSN 0020-9910,1432-1297, Document, Link, MathReview (Jan Nekovář) Cited by: §16.3.
  • [BRE97] C. Breuil (1997) Représentations $p$-adiques semi-stables et transversalité de Griffiths. Math. Ann. 307 (2), pp. 191–224. External Links: ISSN 0025-5831, Document, Link, MathReview (Abdellah Mokrane) Cited by: §11.3, §11, Remark 7.2.
  • [BRI08] O. Brinon (2008) Représentations $p$-adiques cristallines et de de Rham dans le cas relatif. Mém. Soc. Math. Fr. (N.S.) (112), pp. vi+159. External Links: ISSN 0249-633X, ISBN 978-2-85629-250-1, Document, Link, MathReview (Laurent N. Berger) Cited by: §10.
  • [DLL+02] H. Diao, K. Lan, R. Liu, and X. Zhu ([2023] ©2023) Logarithmic adic spaces: some foundational results. pp. 65–182. External Links: ISBN 978-3-031-21549-0; 978-3-031-21550-6, Document, Link, MathReview (Huy Dang) Cited by: §10, §10, §13, §8.
  • [DLL+23] H. Diao, K. Lan, R. Liu, and X. Zhu (2023) Logarithmic Riemann-Hilbert correspondences for rigid varieties. J. Amer. Math. Soc. 36 (2), pp. 483–562. External Links: ISSN 0894-0347,1088-6834, Document, Link, MathReview (Fumio Hazama) Cited by: §10, §11.1, Corollary 11.1, §12, Proposition 12.1, §7, §7, §8, §8, §8, Remark 8.1, Remark 8.4, §8, §9.
  • [EY20] V. Ertl and K. Yamada (2020) Poincaré duality for rigid analytic Hyodo–Kato theory. arXiv. External Links: Document, Link Cited by: §11.6, Lemma 15.4, §15, §16.5, §6.1, §6.
  • [FC90] G. Faltings and C. Chai (1990) Degeneration of abelian varieties. Ergebnisse der Mathematik und ihrer Grenzgebiete (3) [Results in Mathematics and Related Areas (3)], Vol. 22, Springer-Verlag, Berlin. Note: With an appendix by David Mumford External Links: ISBN 3-540-52015-5, Document, Link, MathReview (Min Ho Lee) Cited by: §8.
  • [ILL02] L. Illusie (2002) An overview of the work of K. Fujiwara, K. Kato, and C. Nakayama on logarithmic étale cohomology. Astérisque (279), pp. 271–322. Note: Cohomologies $p$-adiques et applications arithmétiques, II External Links: ISSN 0303-1179,2492-5926, MathReview (Mark Kisin) Cited by: §13.
  • [LLZ23] K. Lan, R. Liu, and X. Zhu (2023) De Rham comparison and Poincaré duality for rigid varieties. Peking Math. J. 6 (1), pp. 143–216. External Links: ISSN 2096-6075,2524-7182, Document, Link, MathReview (Arnaud Mayeux) Cited by: §1.2, §1.2, §10, §10, Proposition 10.1, §11.6, Proposition 12.1, §13, §13, §13, §15, §16.6, §7, Remark 7.1, §7, §7, §9.
  • [MAD19] K. Madapusi Pera (2019) Toroidal compactifications of integral models of Shimura varieties of Hodge type. Ann. Sci. Éc. Norm. Supér. (4) 52 (2), pp. 393–514. External Links: ISSN 0012-9593,1873-2151, Document, Link, MathReview (Brandon Levin) Cited by: §8, §8.
  • [OGU18] A. Ogus (2018) Lectures on logarithmic algebraic geometry. Cambridge Studies in Advanced Mathematics, Vol. 178, Cambridge University Press, Cambridge. External Links: ISBN 978-1-107-18773-3 Cited by: Definition 1.
  • [SW13] P. Scholze and J. Weinstein (2013) Moduli of $p$-divisible groups. Camb. J. Math. 1 (2), pp. 145–237. External Links: ISSN 2168-0930,2168-0949, Document, Link, MathReview (Rui Miguel Saramago) Cited by: §8.
  • [SHI00] A. Shiho (2000) Crystalline fundamental groups. I. Isocrystals on log crystalline site and log convergent site. J. Math. Sci. Univ. Tokyo 7 (4), pp. 509–656. External Links: ISSN 1340-5705, MathReview (Bruno Chiarellotto) Cited by: §2.
  • [SHI02] A. Shiho (2002) Crystalline fundamental groups. II. Log convergent cohomology and rigid cohomology. J. Math. Sci. Univ. Tokyo 9 (1), pp. 1–163. External Links: ISSN 1340-5705, MathReview (Bruno Chiarellotto) Cited by: §2, Proposition 2.2, §2, Remark 3.2, §4.1, §4.5, §4.5, §4.5.
  • [TT19] F. Tan and J. Tong (2019) Crystalline comparison isomorphisms in $p$-adic Hodge theory: the absolutely unramified case. Algebra Number Theory 13 (7), pp. 1509–1581. External Links: ISSN 1937-0652, Document, Link, MathReview (Sinan Ünver) Cited by: §7, §7, §8.